# Interpretation of the Boundary Current Synchronization as a Maxwell’s Demon

### v1

@@ 3.2 Regression analysis
Table 1 shows the estimation results of the coefficients of the dynamical system, where the brackets show the 95% confidence intervals by the bootstrap method. The relaxation coefficients $r_{G}$ and $r_{K}$ are positive, and their inverses give the relaxation time scales. The time scale is about $3$ months for the Gulf Stream and the Kuroshio Current. The estimates of the interaction coefficients $c_{G\leftarrow K}$ and $c_{K\leftarrow G}$ are positive. The confidence intervals suggest that $c_{G\leftarrow K}$ based on the OISST data may be negative. The impact of this negative value is discussed in Appendix C. The variances of the noise, $D_{G}$ and $D_{K}$, are of similar magnitude. We confirmed that the uncertainty quantification is not strongly dependent on the choice of method, as similar results were obtained using a parametric approach based on normality (not shown).
@@ 3.4 Reproduction of the BCS by the dynamical system
Figure 4 shows the resultant lag correlations. Here, the 95% confidence interval for each lag correlation coefficient was estimated using the $2000$ sets of resampled time series obtained by the moving block bootstrap method. The solid lines indicate that the correlation coefficient takes significant values larger than zero, which supports the existence of the synchronicity. The longer interval of significant correlations for the GFDL-CM4C192 data is likely due to its longer data range compared to that of the OISST data. Indeed, the obtained lag correlation coefficients are very close to the results of Kohyama et al. (2021) (see their Fig. 1D). Therefore, we can conclude that the dynamical system serves as an appropriate model for describing BCS.
@@ 4.1 Application of the second law of information thermodynamics
We apply the second law of information thermodynamics (Horowitz and Esposito, 2014; Loos and Klapp, 2020) to the Gulf Stream $T_{G}$ and the Kuroshio Current $T_{K}$ to clarify their asymmetric roles in the BCS. We assume a statistically steady state in which the probability distribution of $T_{G}$ and $T_{K}$ does not vary over time. The time-series analysis supports the validity of the steady-state assumption (Section 33.3).
@@
where $\dot{\sigma}_{G}$ and $\dot{\sigma}_{K}$ represent the production rates of total entropy for the Gulf Stream $T_{G}$ and the Kuroshio Current $T_{K}$, respectively, and $\dot{I}_{G\leftarrow K}$ and $\dot{I}_{K\leftarrow G}$ represent the information flow to the Gulf Stream and to the Kuroshio Current, respectively. Equations (21), (22), and (23) correspond to Eqs. (14), (15), and (13), respectively. The sum of the two information flows is zero in Eq. (23) because the time derivative of mutual information is zero due to the steady-state assumption (Appendix A). Moreover, under this assumption, the production rates of total entropy correspond to $\Delta Q$ in Eq. (9) due to the Shannon entropies being zero. The interpretation of $\Delta Q$ for the dynamical system is discussed in Subsection .5 in Appendix A. Adding Eqs. (21) and (22) yields $\dot{\sigma}_{G}+\dot{\sigma}_{K}\geq 0$, i.e., the total entropy production for the entire system is non-negative, indicating the consistency with the second law of thermodynamics. Considering each subsystem, $\dot{\sigma}_{G}$ and $\dot{\sigma}_{K}$ are lower bounded by $\dot{I}_{G\leftarrow K}$ and $\dot{I}_{K\leftarrow G}$, respectively, one of which can be negative. By using information thermodynamics, we investigate the asymmetry between the dynamics of $T_{G}$ and $T_{K}$ from the perspective of entropy production and information transfer.
@@ 4.2 Analysis of regime diagrams for the BCS
Figures 5 and 6 indicate that the dynamical system is a Maxwell’s demon system, where the Gulf Stream is a “Particle” and the Kuroshio Current is a “Demon.” The entropy production rate and information flow of the Gulf Stream, $\dot{\sigma}_{G}$ and $\dot{I}_{G\leftarrow K}$, are negative (Figs. 5a, 5b, 6a, and 6b), whereas those of the Kuroshio Current, $\dot{\sigma}_{K}$ and $\dot{I}_{K\leftarrow G}$, are positive (Figs. 5d, 5e, 6d, and 6e). According to the discussions on Maxwell’s demon (Sections 22.3 and 22.4), the Kuroshio Current as the “Demon” gains mutual information by measuring the Gulf Stream SST, which is reflected in the positive information flow ($\dot{I}_{K\leftarrow G}>0$). Furthermore, the Kuroshio Current utilizes its own SST as a memory and exhibits a positive entropy production rate ($\dot{\sigma}_{K}>0$), which is attributed to the memory update process associated with the measurement. On the other hand, the Gulf Stream as the “Particle” loses mutual information due to the feedback control by the Kuroshio Current, which is reflected in the negative information flow ($\dot{I}_{G\leftarrow K}<0$). In addition, the Gulf Stream SST varies by rectifying noise and shows a negative entropy production rate ($\dot{\sigma}_{K}<0$). The term “rectify” means to extract coherent directional output from random noise fluctuations.
We further discuss the rectification of noise. Figures 5c, 5f, 6c, and 6f show that the SST magnitudes, $\sqrt{\langle T_{G}^{2}\rangle}$ and $\sqrt{\langle T_{K}^{2}\rangle}$, are increased compared to the case without interactions. As discussed in Section 33.2, without noise, the dynamical system would decay to zero without oscillation. This result suggests that the driving source is noise. Moreover, in the absence of interactions, namely $c_{G\leftarrow K}=c_{K\leftarrow G}=0$, the SSTs of the Gulf Stream and the Kuroshio Current would evolve independently, and BCS would not occur. Compared to the SST magnitudes in this case, they are increased due to the interactions. Within the framework of the dynamical system, the presence of both interactions and noises is essential for BCS.
@@
The negative $\dot{\sigma}_{G}$ suggests that the action from the Kuroshio Current on the Gulf Stream, $c_{G\leftarrow K}T_{K}$, does not directly contribute to the increase in $T_{G}$. Indeed, this action $c_{G\leftarrow K}T_{K}$ is negatively correlated with the tendency of the Gulf Stream SST, $c_{G\leftarrow K}\langle T_{K}\,{\rm d}T_{G}/{\rm d}t\rangle=D_{G}\dot{\sigma}_{G}<0$. This result is due to the relaxation. Without relaxation, the governing equation (16) would be ${\rm d}T_{G}/{\rm d}t=c_{G\leftarrow K}T_{K}$, and only the first term in Eq. (31) would remain, leading to a positive $\dot{\sigma}_{G}$. The mechanism for the increase in $T_{G}$ is then described as follows. For example, when $T_{G}$ is positive, it tends to decrease to zero due to relaxation, but this decrease is mitigated by $c_{G\leftarrow K}T_{K}$. In other words, the action from the Kuroshio Current suppresses the relaxation of the Gulf Stream. Consequently, the remaining term $\xi_{G}$ in Eq. (16), namely the noise, becomes more effective and increase $T_{G}$. This suppression of relaxation to zero corresponds to the feedback control of Maxwell’s demon that inserts a partition (Fig. 7).
@@
The positive $\dot{\sigma}_{K}$ suggests that the action from the Gulf Stream on the Kuroshio Current, $c_{K\leftarrow G}T_{G}$, directly contributes to the increase in $T_{K}$. Similar to the Gulf Stream, there is an effect of suppressing relaxation; however, its contribution is relatively small and $\dot{\sigma}_{K}$ is positive. The amplification of $T_{K}$ occurs by following the changes in $T_{G}$ due to the action $c_{K\leftarrow G}T_{G}$ from the Gulf Stream. According to the framework of Maxwell’s demon, the Kuroshio Current (i.e., the “Demon”) uses its SST as memory and the SST fluctuations are increased by updating the memory value $T_{K}$ based on the measurement outcome of the Gulf Stream SST.
@@
4.3 Interpretation of BCS as a Maxwell’s demon system
@@
Figure 9: Schematic of the interpretation of the BCS as a Maxwell’s demon system. The Kuroshio Current plays the role of the “Demon” and the Gulf Stream plays the role of the “Particle.” See Section 44.3 for details.
@@ 5 Conclusions
This study has introduced a bivariate linear dynamical system that describes the BCS and analyzed it using a theory of information thermodynamics (Horowitz and Esposito, 2014; Loos and Klapp, 2020). This system can be interpreted as a Maxwell’s demon system, with the Gulf Stream playing the role of the “Particle” and the Kuroshio Current playing the role of the “Demon.” Information thermodynamics clarifies the asymmetric roles of the Gulf Stream and the Kuroshio Current from the perspective of entropy production and information transfer (Section 44.3). Our results suggest that random noise fluctuations can be autonomously converted into coherent, directional signals, or in other words, into essential variations, in the climate system.
For future research, we propose applying information thermodynamics to more realistic models that describe the BCS (e.g., Gallego and Cessi, 2001; Kohyama et al., 2021). While the current study demonstrates that the Gulf Stream can be interpreted as the “Particle” and the Kuroshio Current as the “Demon,” these roles might be reversed under certain conditions, such as due to temporal changes in interaction coefficients. Moreover, it is unclear how the interactions, such as the interference of the relaxation, are realized in the real climate system. These points will be revealed by applying the theory to realistic models.
@@ Acknowledgements.
The information flow $\dot{I}_{i\leftarrow}$ represents the mutual information gained by the subsystem $x_{i}$ per unit time. We further discuss information flow in Subsection .3 in Appendix A. In Eq. (45), the information flow $\dot{I}_{\zeta\leftarrow}$ (i.e., $i=\zeta$) was substituted.
@@
$f(\mathbf{x}(t))\circ\frac{{\rm d}z}{{\rm d}t}$ $=f(\mathbf{x}(t))\circ\left(-a_{\zeta}+D_{\zeta}\frac{{\rm d}W_{\zeta}}{{\rm d}t}\right),$ (54)
$=-f(\mathbf{x}(t))a_{\zeta}+D_{\zeta}f(\bm{x}(t))\circ\frac{{\rm d}W_{\zeta}}{{\rm d}t},$ (55)
@@
The sum of the information gained by the subsystem $x_{i}$ per unit time (i.e., the information flow $\dot{I}_{i\leftarrow}$) over all $i$ becomes the temporal change of the mutual information of the entire system. The information flow $\dot{I}_{i\leftarrow}$ is interpreted as the sum of the inflow and outflow of information from all the other subsystems to $x_{i}$.
@@
The present study has analyzed the dynamical system [Eqs. (16) and (17)] using the second law of information thermodynamics as a mathematical tool (Sections 22.4 and 4). It is worth noting that mathematically deriving the second law as a constraint is different from obtaining a meaningful physical insight from it. We have regarded $\dot{\sigma}_{G}$ ($=\dot{Q}_{G}/D_{G}$) and $\dot{\sigma}_{K}$ ($=\dot{Q}_{K}/D_{K}$) as indicators of the contribution of interactions to the temporal tendencies of the SSTs. This interpretation does not rely on the interpretation of $\dot{Q}_{i}$ as heat. As a similar example, Ito and Sagawa (2015) considered $\dot{Q}_{i}$ as the robustness of signal transmission in a cell and analyzed the processing efficiency of cell signaling using information thermodynamics. The application of information thermodynamics in this manner might yield interesting results for other atmospheric and oceanic systems. Investigating further applications is an important topic for future research in atmospheric and oceanic sciences.
@@
where Eqs. (98) and (99) are substituted. Equations (104) and (106) are the same as (24) and (25), respectively. Equation (29) is obtained using Eqs. (100), (104), and (106).
@@
$I_{i\leftarrow}$ $=-\int{\rm d}\mathbf{x}\;\frac{\partial J_{i}}{\partial x_{i}}\ln p_{\rm tot},$ (107)
@@
$I_{G\leftarrow K}$ $=-(\bm{\mathsf{\Sigma}}^{-1})_{KG}\left\langle T_{K}\circ\frac{{\rm d}T_{G}}{{\rm d}t}\right\rangle,$ (110)
$I_{K\leftarrow G}$ $=-(\bm{\mathsf{\Sigma}}^{-1})_{GK}\left\langle T_{G}\circ\frac{{\rm d}T_{K}}{{\rm d}t}\right\rangle,$ (111)
@@
The information flows and entropy production rates become zero when $c_{G\leftarrow K}D_{K}=c_{K\leftarrow G}D_{G}$ (Loos and Klapp, 2020). This fact can be directly verified by substituting Eqs. (91)–(93), (101), and (102) into (24)–(27). A simpler method to show this is through the application of the second law of information thermodynamics (see Appendix D). In the case of $c_{G\leftarrow K}D_{K}=c_{K\leftarrow G}D_{G}$, the two subsystems, namely the Gulf Stream and the Kuroshio Current, become symmetric. Mathematically, the detailed balance is satisfied and the probability flux $\mathbf{J}$ is zero (Loos and Klapp, 2020).
@@
We quantified the uncertainty of the estimated coefficients using the moving block bootstrap method (Mudelsee, 2014) in Section 33.2, where the number of resampled time series was set to $2000$. We discuss here the histograms of the entropy production rates and information flows obtained from these coefficients.
@@
In the region $c_{G\leftarrow K}<0$, the interpretation of Maxwell’s demon is not possible because of the same signs of $\dot{\sigma}_{G}$ and $\dot{\sigma}_{K}$, but an interesting phenomenon still occurs in terms of information thermodynamics, which is called feedback cooling (Horowitz and Sandberg, 2014; Loos and Klapp, 2020). For the interpretation of Maxwell’s demon, the signs of the entropy production rates need to be opposite to each other (Sections 22.3 and 22.4). As for the Kuroshio Current, the entropy production rate and information flow are positive, as in the case of Maxwell’s demon. Thus, even in the scenario of feedback cooling, there is no change in the interpretation for the Kuroshio Current. The difference arises in the Gulf Stream, which has a positive entropy production rate ($\dot{\sigma}_{G}>0$). The Gulf Stream still exhibits a negative information flow, which indicates that the Gulf Stream is interpreted as being controlled by the Kuroshio Current, like in Maxwell’s demon. However, this control is done by negative feedback and reduces the effect of noise (Horowitz and Sandberg, 2014; Loos and Klapp, 2020). As a result, the SST magnitude of the Gulf Stream is reduced (Fig. 5c). This reduction is considered as being achieved by discarding heat to the environment, leading to the positive entropy production rate ($\dot{\sigma}_{G}>0$). Even in this parameter region, the correlation between $T_{G}$ and $T_{K}$ is positive (Fig. 8), suggesting that the BCS occurs. However, the interpretation is slightly different. The synchronization is achieved by the Kuroshio Current, which reduces the SST fluctuations of the Gulf Stream caused by atmospheric and oceanic noise.

### v2

@@ 3.1 Governing equations
This type of a system that consists only of oceanic variables (e.g., SSTs) can be derived from a more complicated setting that include both oceanic and atmospheric variables. Indeed, Gallego and Cessi (2001) started with an atmosphere-ocean coupling model and then derived idealized equations that consist only of Atlantic and Pacific variables by representing the atmospheric effects in terms of the oceanic variables. Our dynamical system is interpreted as a simplified model derived through these procedures. The cross terms, $c_{G\leftarrow K}T_{K}$ and $c_{K\leftarrow G}T_{G}$, may include inter-basin couplings via the atmosphere.
@@ 3.2 Regression analysis
Table 1 shows the estimation results of the coefficients of the dynamical system, where the brackets show the 95% confidence intervals by the bootstrap method. The relaxation coefficients $r_{G}$ and $r_{K}$ are positive, and their inverses give the relaxation time scales. The time scale is about $3$ months for the Gulf Stream and the Kuroshio Current. The estimates of the interaction coefficients $c_{G\leftarrow K}$ and $c_{K\leftarrow G}$ are positive. The confidence intervals suggest that $c_{G\leftarrow K}$ based on the OISST data may be negative. The impact of this negative value is discussed in Appendix C. The variances of the noise $D_{G}$ and $D_{K}$ are of similar magnitude. We confirmed that the uncertainty quantification is not strongly dependent on the choice of method, as similar results were obtained using a parametric approach based on normality (not shown).
@@ 3.3 Validation of the assumptions for the dynamical system
We finally clarify the relationship between noise and temperature. In non-equilibrium physics, the variance of stochastic noise is generally proportional to the temperature of an isothermal environment; this noise arises from collisions of small molecules with the larger target particle (e.g., Sekimoto, 2010). In our case, since the noises $\xi_{G}$ and $\xi_{K}$ represent atmospheric and oceanic fluctuations, their variances $D_{G}$ and $D_{K}$ are interpreted as the strengths of these small-scale disturbances. Thus, isothermal environments correspond to the disturbance fields, and “temperature” in non-equilibrium physics corresponds to $D_{G}$ and $D_{K}$. These effective temperatures should be distinguished from the SST anomalies $T_{G}$ and $T_{K}$ at longer time scales.
@@ 3.4 Reproduction of the BCS by the dynamical system
Figure 4 shows the resultant lag correlations. Here, the 95% confidence interval for each lag correlation coefficient was estimated using the $2000$ sets of resampled time series obtained by the moving block bootstrap method. The solid lines indicate that the correlation coefficient takes significant values larger than zero, which supports the existence of the synchronicity. The longer interval of significant correlations for the GFDL-CM4C192 data is likely due to its longer data range compared to that of the OISST data. The Gulf Stream slightly leads the Kuroshio Current by 1 and 0.2 months for the OISST and GFDL-CM4C192 data, respectively. The significance of these small leads is difficult to assess given the short data periods. The obtained lag correlation coefficients are very close to the results of Kohyama et al. (2021) (see their Fig. 1D). Therefore, we can conclude that the dynamical system serves as an appropriate model for describing BCS.
@@ 4.1 Application of the second law of information thermodynamics
We apply the second law of information thermodynamics (Horowitz and Esposito, 2014; Loos and Klapp, 2020) to the Gulf Stream $T_{G}$ and the Kuroshio Current $T_{K}$ to clarify their asymmetric roles in the BCS. We assume a statistical steady state in which the probability distribution of $T_{G}$ and $T_{K}$ does not vary over time. The time-series analysis supports the validity of the steady-state assumption (Section 33.3).
@@
where $\dot{\sigma}_{G}$ and $\dot{\sigma}_{K}$ represent the production rates of total entropy for the Gulf Stream $T_{G}$ and the Kuroshio Current $T_{K}$, respectively, and $\dot{I}_{G\leftarrow K}$ and $\dot{I}_{K\leftarrow G}$ represent the information flow to the Gulf Stream and to the Kuroshio Current, respectively. Equations (21), (22), and (23) correspond to Eqs. (14), (15), and (13), respectively. The sum of the two information flows is zero in Eq. (23) because the time derivative of mutual information is zero due to the steady-state assumption (Appendix A). Moreover, under this assumption, the production rates of total entropy correspond to $\Delta Q$ in Eq. (9) because the change of Shannon entropies is zero. The interpretation of $\Delta Q$ for the dynamical system is discussed in Subsection .5 in Appendix A. Adding Eqs. (21) and (22) yields $\dot{\sigma}_{G}+\dot{\sigma}_{K}\geq 0$, i.e., the total entropy production for the entire system is non-negative, indicating the consistency with the second law of thermodynamics. Considering each subsystem, $\dot{\sigma}_{G}$ and $\dot{\sigma}_{K}$ are lower bounded by $\dot{I}_{G\leftarrow K}$ and $\dot{I}_{K\leftarrow G}$, respectively, one of which can be negative. By using information thermodynamics, we investigate the asymmetry between the dynamics of $T_{G}$ and $T_{K}$ from the perspective of entropy production and information transfer.
@@ 4.2 Analysis of regime diagrams for the BCS
Figures 5 and 6 indicate that the dynamical system is a Maxwell’s demon system, where the Gulf Stream is a “Particle” and the Kuroshio Current is a “Demon.” The entropy production rate and information flow of the Gulf Stream, $\dot{\sigma}_{G}$ and $\dot{I}_{G\leftarrow K}$, are negative (Figs. 5a, 5b, 6a, and 6b), whereas those of the Kuroshio Current, $\dot{\sigma}_{K}$ and $\dot{I}_{K\leftarrow G}$, are positive (Figs. 5d, 5e, 6d, and 6e). According to the discussions on Maxwell’s demon systems (Sections 22.3 and 22.4), the Kuroshio Current as the “Demon” tends to establish the correlation with the Gulf Stream SST by measuring it, which is reflected in the positive information flow ($\dot{I}_{K\leftarrow G}>0$). Furthermore, the Kuroshio Current utilizes its own SST as a memory and exhibits a positive entropy production rate ($\dot{\sigma}_{K}>0$), which is attributed to the memory update process associated with the measurement. On the other hand, the Gulf Stream as the “Particle” tends to decrease the correlation due to the feedback control by the Kuroshio Current, which is reflected in the negative information flow ($\dot{I}_{G\leftarrow K}<0$). In addition, the Gulf Stream SST varies by rectifying noise and shows a negative entropy production rate ($\dot{\sigma}_{G}<0$). The term “rectify” means to extract coherent directional output from random noise fluctuations.
We further discuss the rectification of noise. Figures 5c, 5f, 6c, and 6f show that the SST magnitudes $\sqrt{\langle T_{G}^{2}\rangle}$ and $\sqrt{\langle T_{K}^{2}\rangle}$ are increased compared to the case without interactions. As discussed in Section 33.2, without noise, the dynamical system would decay to zero without oscillation. This result suggests that the driving source is noise. Moreover, in the absence of interactions, namely $c_{G\leftarrow K}=c_{K\leftarrow G}=0$, the SSTs of the Gulf Stream and the Kuroshio Current would evolve independently, and BCS would not occur. Compared to the SST magnitudes in this case, they are increased due to the interactions. Within the framework of the dynamical system, the presence of both interactions and noises is essential for BCS.
@@
The negative $\dot{\sigma}_{G}$ suggests that the action from the Kuroshio Current on the Gulf Stream, $c_{G\leftarrow K}T_{K}$, does not directly contribute to the increase in $T_{G}$. Indeed, this action $c_{G\leftarrow K}T_{K}$ is negatively correlated with the tendency of the Gulf Stream SST, $c_{G\leftarrow K}\langle T_{K}\,{\rm d}T_{G}/{\rm d}t\rangle=D_{G}\dot{\sigma}_{G}<0$. This result is due to the relaxation. Without relaxation, the governing equation (16) would be ${\rm d}T_{G}/{\rm d}t=c_{G\leftarrow K}T_{K}$, and only the first term in Eq. (31) would remain, leading to a positive $\dot{\sigma}_{G}$. The mechanism for the increase in $T_{G}$ is then described as follows. For example, when $T_{G}$ is positive, it tends to decrease to zero due to relaxation, but this decrease is mitigated by $c_{G\leftarrow K}T_{K}$. In other words, the action from the Kuroshio Current suppresses the relaxation of the Gulf Stream. Consequently, the remaining term $\xi_{G}$ in Eq. (16), namely the noise, becomes more effective and increase $T_{G}$. This suppression of relaxation to zero corresponds to the feedback control that inserts a partition (Fig. 7).
@@
The positive $\dot{\sigma}_{K}$ suggests that the action from the Gulf Stream on the Kuroshio Current, $c_{K\leftarrow G}T_{G}$, directly contributes to the increase in $T_{K}$. Similar to the Gulf Stream, there is an effect of suppressing relaxation; however, its contribution is relatively small and $\dot{\sigma}_{K}$ is positive. The amplification of $T_{K}$ occurs by following the changes in $T_{G}$ due to the action $c_{K\leftarrow G}T_{G}$ from the Gulf Stream. According to the framework of Maxwell’s demon systems, the Kuroshio Current (i.e., the “Demon”) uses its SST as memory and the SST fluctuations are increased by updating the memory value $T_{K}$ based on the measurement outcome of the Gulf Stream SST.
These analyses of $T_{G}$ and $T_{K}$ suggest that information-to-energy conversion (Toyabe et al., 2010) occurs in a subsystem (i.e., the Gulf Stream), even though the entire system is dissipative, as in other Maxwell demon systems. Generally, stochastic noise fluctuates a system back and forth, so the net energy or work is difficult to extract. However, using the information from measurements, fluctuations can be converted into directional movements in a subsystem (Toyabe et al., 2010), corresponding to the negative $\dot{\sigma}_{G}$ or increased $T_{G}$ amplitude in our system. Such conversion cannot occur for the entire dissipative system (i.e., no net positive conversion), as reflected in the positive value of $\dot{\sigma}_{G}+\dot{\sigma}_{K}$.
Synchronicity is essential for this conversion. For instance, in Fig. 7, the partition needs to be inserted just behind the “Particle.” This insertion is based on the memory value of the “Demon,” as in the Szilard engine (Fig. 1), where the memory value becomes equal to the particle position after the measurement. Although it is difficult to directly compare our dynamical system to the Szilard engine because our system is autonomous (Section 22.4), the memory value $X$ and particle position $Y$ correspond to $T_{K}$ and $T_{G}$, respectively. Both subsystems $T_{K}$ and $T_{G}$ are influenced by the small-scale disturbance fields (Section 33.3), corresponding to isothermal environments in the Szilard engine. The “Demon” ($T_{K}$) follows a change in the “Particle” ($T_{G}$), thereby enabling feedback control to convert fluctuations. The $T_{K}$ amplitude increases due to this following (i.e., its continual updates based on $T_{G}$), although $T_{K}$ is dissipative (i.e., $\dot{\sigma}_{K}>0$). In the Szilard engine, this dissipation corresponds to the work required to initialize the memory.
@@
4.3 Interpretation of BCS as a Maxwell’s demon system: stochastic synchronization
@@
Figure 9: Schematic of the interpretation of the BCS as a Maxwell’s demon system (i.e., the stochastic synchronization). The Kuroshio Current plays the role of the “Demon” and the Gulf Stream plays the role of the “Particle.” See Section 44.3 for details.
In this regard, Yamagami et al. (2024) recently examined the asymmetric roles between the Gulf Stream and Kuroshio Current using a coupled atmosphere-ocean general circulation model through “pacemaker” experiments. In these experiments, SST variations only in a western boundary current region are strongly relaxed toward (i.e., keep “pace” with) a control run, so that the responses of the other climate systems to the western boundary current variability are isolated. They showed that, when the Gulf Stream acts as a pacemaker, BCS is reproduced. The heat release from the Gulf Stream modulates the atmospheric annular mode, exciting oceanic Rossby waves in the Pacific and influencing the SST in the Kuroshio Current region. This mechanism serves as a plausible physical process that realizes the role of the “Particle” in the context of the Maxwell’s demon system, i.e., the Gulf Stream forces the SST of the Kuroshio Current to be in phase. Conversely, they also showed that, when the Kuroshio Current acts as a pacemaker, the BCS is not reproduced. This result is also consistent with our interpretation that, unlike the Gulf Stream, the Kuroshio Current does not force the SST of the Gulf Stream by itself. Rather, by responding to the Gulf Stream’s forcing, the Kuroshio Current returns a feedback to interfere with the Gulf Stream SST relaxation toward its climatology. The absence of BCS in the Kuroshio pacemaker experiment can be understood that, because its SST varies independently from the state of the Gulf Stream, the Kuroshio Current cannot reproduce an optimal feedback in an appropriate timing. Our future work is to reveal what kind of physical processes regarding the Kuroshio Current realizes the role of the “Demon” in the real world.
Our proposed mechanism can be realized in other stochastic dissipative systems with positive feedback. The dynamical system investigated here, Eqs. (16) and (17), is known as the Langevin system, one of the simplest and most common models with dissipation and stochastic forcing (e.g., Sekimoto, 2010). Indeed, Langevin systems have been employed as simple models for climate phenomena (e.g., Dijkstra, 2013). The key ingredient in our study is the positive feedback; that is, both interaction coefficients $c_{K\leftarrow G}$ and $c_{G\leftarrow K}$ are positive. Due to this positive feedback, atmospheric or oceanic fluctuations are converted into directional variations in a subsystem via synchronization (i.e., information-to-energy conversion; Toyabe et al. (2010)), although such conversion cannot be realized for the entire dissipative system. The wide applicability of Langevin systems suggests that the proposed mechanism of synchronization (Fig. 9), “stochastic synchronization,” may describe other climate phenomena. Such applications remain an important topic for future work.
@@ 5 Conclusions
This study has introduced a bivariate linear dynamical system that describes the BCS and analyzed it using a theory of information thermodynamics (Horowitz and Esposito, 2014; Loos and Klapp, 2020). This system can be interpreted as a Maxwell’s demon system, with the Gulf Stream playing the role of the “Particle” and the Kuroshio Current playing the role of the “Demon.” Information thermodynamics clarifies the asymmetric roles of the Gulf Stream and the Kuroshio Current from the perspective of entropy production and information transfer (Section 44.3). Our results suggest that random noise fluctuations can be autonomously converted into coherent, directional variations (i.e., information-to-energy conversion), through synchronization between the Gulf Stream and the Kuroshio Current. The proposed mechanism, named “stochastic synchronization,” may describe other climate phenomena because our analysis was based on a widely applicable Langevin system.
For future research, we propose applying information thermodynamics to more realistic models that describe the BCS (e.g., Gallego and Cessi, 2001; Kohyama et al., 2021). While the current study demonstrates that the Gulf Stream can be interpreted as the “Particle” and the Kuroshio Current as the “Demon,” what determines these roles remains unclear, and these roles might be reversed under certain conditions, such as due to temporal changes in interaction coefficients. Moreover, it is unclear how the interactions, such as the interference of the relaxation, are realized in the real climate system. These points will be revealed by applying the theory to realistic models.
@@ Acknowledgements.
The information flow $\dot{I}_{i\leftarrow}$ represents changes in the mutual information by the subsystem $x_{i}$. We further discuss information flow in Subsection .3 in Appendix A. In Eq. (45), the information flow $\dot{I}_{\zeta\leftarrow}$ (i.e., $i=\zeta$) was substituted.
@@
$f(\mathbf{x}(t))\circ\frac{{\rm d}z}{{\rm d}t}$ $=f(\mathbf{x}(t))\circ\left(-a_{\zeta}+\sqrt{2D_{\zeta}}\frac{{\rm d}W_{\zeta}}{{\rm d}t}\right),$ (54)
$=-f(\mathbf{x}(t))a_{\zeta}+\sqrt{2D_{\zeta}}f(\bm{x}(t))\circ\frac{{\rm d}W_{\zeta}}{{\rm d}t},$ (55)
@@
The sum of changes in the mutual information by the subsystem $x_{i}$ (i.e., the information flow $\dot{I}_{i\leftarrow}$) over all $i$ becomes the temporal change of the mutual information of the entire system. The information flow $\dot{I}_{i\leftarrow}$ is interpreted as the sum of the inflow and outflow of information from all the other subsystems to $x_{i}$.
@@
The present study has analyzed the dynamical system [Eqs. (16) and (17)] using the second law of information thermodynamics as a mathematical tool (Section 4). Although the second law is always derived for SDEs as shown above, mathematically deriving the second law is different from obtaining a meaningful physical insight from it. We have regarded $\dot{\sigma}_{G}$ ($=\dot{Q}_{G}/D_{G}$) and $\dot{\sigma}_{K}$ ($=\dot{Q}_{K}/D_{K}$) as indicators of the contribution of interactions to the temporal tendencies of the SSTs. This interpretation does not rely on the interpretation of $\dot{Q}_{i}$ as heat. As a similar example, Ito and Sagawa (2015) considered $\dot{Q}_{i}$ as the robustness of signal transmission in a cell (i.e., not as heat) and analyzed the processing efficiency of cell signaling using information thermodynamics. The application of information thermodynamics in this manner might yield interesting results for other atmospheric and oceanic systems. Investigating further applications is an important topic for future research in atmospheric and oceanic sciences.
@@
where Eqs. (98) and (99) are substituted. Equations (104) and (106) are the same as Eqs. (24) and (25), respectively. Equation (29) is obtained using Eqs. (100), (104), and (106).
@@
$\dot{I}_{i\leftarrow}$ $=-\int{\rm d}\mathbf{x}\;\frac{\partial J_{i}}{\partial x_{i}}\ln p_{\rm tot},$ (107)
@@
$\dot{I}_{G\leftarrow K}$ $=-(\bm{\mathsf{\Sigma}}^{-1})_{KG}\left\langle T_{K}\circ\frac{{\rm d}T_{G}}{{\rm d}t}\right\rangle,$ (110)
$\dot{I}_{K\leftarrow G}$ $=-(\bm{\mathsf{\Sigma}}^{-1})_{GK}\left\langle T_{G}\circ\frac{{\rm d}T_{K}}{{\rm d}t}\right\rangle,$ (111)
@@
The information flows and entropy production rates become zero when $c_{G\leftarrow K}D_{K}=c_{K\leftarrow G}D_{G}$ (Loos and Klapp, 2020). This fact can be directly verified by substituting Eqs. (91)–(93), (101), and (102) into Eqs. (24)–(27). A simpler method to show this is through the application of the second law of information thermodynamics (see Appendix D). In the case of $c_{G\leftarrow K}D_{K}=c_{K\leftarrow G}D_{G}$, the two subsystems, namely the Gulf Stream and the Kuroshio Current, become symmetric. Mathematically, the detailed balance is satisfied and the probability flux $\mathbf{J}$ is zero (Loos and Klapp, 2020).
@@
We quantified the uncertainty of the estimated coefficients using the moving block bootstrap method (Mudelsee, 2014) in Section 33.2, where the number of resampled time series was set to $2000$. We discuss here the histograms of the entropy production rates and information flows obtained from the coefficients estimated by the bootstrap method.
@@
In the region $c_{G\leftarrow K}<0$, the interpretation of Maxwell’s demon systems is not possible because of the same signs of $\dot{\sigma}_{G}$ and $\dot{\sigma}_{K}$, but an interesting phenomenon still occurs in terms of information thermodynamics, which is called feedback cooling (Horowitz and Sandberg, 2014; Loos and Klapp, 2020). For the interpretation of Maxwell’s demon systems, the signs of the entropy production rates need to be opposite to each other (Sections 22.3 and 22.4). As for the Kuroshio Current, the entropy production rate and information flow are positive, as in the case of Maxwell’s demon. Thus, even in the scenario of feedback cooling, there is no change in the interpretation for the Kuroshio Current. The difference arises in the Gulf Stream, which has a positive entropy production rate ($\dot{\sigma}_{G}>0$). The Gulf Stream still exhibits a negative information flow, which indicates that the Gulf Stream is interpreted as being controlled by the Kuroshio Current, like in Maxwell’s demon. However, this control is done by negative feedback and reduces the effect of noise (Horowitz and Sandberg, 2014; Loos and Klapp, 2020). As a result, the SST magnitude of the Gulf Stream is reduced (Fig. 5c). This reduction is considered as being achieved by discarding heat to the environment, leading to the positive entropy production rate ($\dot{\sigma}_{G}>0$). Even in this parameter region, the correlation between $T_{G}$ and $T_{K}$ is positive (Fig. 8), suggesting that the BCS occurs. However, the interpretation is slightly different. The synchronization is achieved by the Kuroshio Current, which reduces the SST fluctuations of the Gulf Stream caused by atmospheric and oceanic noise.